\begin{thebibliography}{99}

\bibitem{Koutras} A. Koutras,
{\it Class. Quant. Grav. }, {\bf 9}, 1573, (1992).

\bibitem{Am-Mah} G. Amery and S. D. Maharaj,
{\it Int. J. Mod. Phys. D}, {\bf 11}, 337, (2002).

\bibitem{ERB} S. B. Edgar, R. Rani, and A. Barbes,
{\it Class. Quant. Grav.}, {\bf 20}, 1929, (2003).

\bibitem{Ya} K. Yano,
{\it Ann. Math.}, {bf 55}, 328, (1952).

\bibitem{Gibbons}G. W. Gibbons, R. H. Rietdijk and J. W. van Holten,
{\it Nucl. Phys. B}, {\bf 404}, 42, (1993).

\bibitem{GiRu1} G. W. Gibbons and P. J. Ruback,
{\it Phys. Lett. B}, {\bf 188}, 226, (1987).

\bibitem{GiRu2} G. W. Gibbons and P. J. Ruback,
{\it Commun. Math. Phys.}, {\bf 115}, 267, (1988).

\bibitem{VaVi1} D. Vaman and M. Visinescu,
{\it Phys. Rev. D}, {\bf 57}, 3790, (1998).

\bibitem{vH} J. W. van Holten,
{\it Phys. Lett. B}, {\bf 342}, 47, (1995).

\bibitem{KT} D. Kastor and J. Traschen,
{\it JHEP}, {\bf 0408}, 45, (2004).

\bibitem{baleanu-baskal} D. Baleanu and S. Baskal,
{\it Int. J. Mod. Phys. A}, {\bf 17}, 3737, (2002).

\bibitem{Ri1} R. H. Rietdijk and J. W. van Holten,
{\it Nucl. Phys. B}, {\bf 472}, 427, (1996).

\bibitem{D1} D. Baleanu and A. Karasu,
{\it Mod. Phys. Lett. A}, {\bf 14}, 2597, (1999).

\bibitem{D2} D. Baleanu and S. Baskal,
{\it Mod. Phys. Lett. A}, {\bf 16}, 135, (2001).

\bibitem{hall} G. S. Hall,
{\it Int. J. Theor. Phys.}, {\bf 26}, 71, (1987).

\bibitem{how} L. Howarth,
Ph.D. thesis, University of Hull, (1999).

\bibitem{HC1} L. Howarth and C. D. Collinson,
{\it Gen. Rel. Grav.}, {\bf 32}, 1845, (2000).

\bibitem{HC2} L. Howarth and C. D. Collinson,
{\it Gen. Rel. Grav}, {\bf 32}, 1767, (2000).

\bibitem{Carter} B. Carter,
{\it Phys. Rev. D}, {\bf 16}, 3395, (1977).

\bibitem{CMcL} B. Carter and R. G. McLenaghan,
{\it Phys. Rev. D}, {\bf 19}, 1093, (1979).

\bibitem{Visi1} M. Visinescu and I. Cotaescu,
{\it Class. Quant. Grav.}, {\bf 21}, 11, (2004).

\bibitem{Visi2} M. Visinescu and I. Cotaescu,
{\it J. Math. Phys.}, {\bf 43}, 2978, (2002).

\bibitem{Visi3} M. Visinescu and I. Cotaescu,
{\it Class. Quant. Grav.}, {\bf 18}, 3383, (2001).

\bibitem{Klishevich} V. V. Klishevich,
{\it Class. Quant. Grav.}, {\bf 17}, 305, (2000).

\bibitem{tanimoto} M. Tanimoto,
{\it Nucl. Phys. B}, {\bf 442}, 549, (1995).

\bibitem{cariglia} M. Cariglia,
{\it Class. Quant. Grav.}, {\bf 21}, 1051, (2004).

\bibitem{Di-Ru} W. Dietz and R. Rudiger,
{\it Proc. R. Soc. Lond. A}, {\bf 375}, 361, (1981).

\bibitem{Hauser-Malhiot} I. Hauser and R. J. Malhiot,
{\it J. Math. Phys.}, {\bf 15}, 816, (1974).

\bibitem{Kimura1} M. Kimura,
{\it Tensors N. S.}, {\bf 30}, 27, (1976).

\end{thebibliography}

\documentclass[twoside,11pt] {article}

\begin{document}

\title{Irreducible Killing Tensors from Third Rank Killing-Yano Tensors }

\author{Florian Catalin Popa \thanks{E-mail:~~~catalin@venus.nipne.ro} \ and Ovidiu Tintareanu-Mircea
\thanks{E-mail:~~~ovidiu@venus.nipne.ro}\\
{\small \it Institute of Space Sciences}\\
{\small \it Atomistilor 409, RO 077125, P.O. Box MG 23}\\
{\small \it Magurele, Bucharest, Romania}}

\date{}
\maketitle

\begin{abstract}

We investigate higher rank Killing-Yano tensors showing that third
rank Killing-Yano tensors are not always trivial objects being
possible to construct irreducible Killing tensors from them.  We
give as an example the Kimura IIC metric were from two rank
Killing-Yano tensors we obtain a reducible Killing tensor and from
third rank Killing-Yano tensors we obtain three Killing tensors,
one reducible and two irreducible.

Key words: Killing tensors, Killing-Yano tensors.
\end{abstract}

\section{Introduction}
The interest in irreducible Killing tensors is due to their
connection with quadratic first integrals, which are not simply
linear combinations of the first integrals associated with Killing
vectors of geodesic motion. Few examples of irreducible Killing
tensors are known explicitly since the direct integration of
Killing equation is not easy. Until now, two indirect methods are
known in the construction of Killing tensors, some of them
expected to be irreducible. In the first one, irreducible Killing
tensors can be considered the "square" of Killing-Yano tensors,
and in the second method Killing tensors can be constructed from
conformal Killing vectors \cite{Koutras, Am-Mah, ERB}.  An
illustration of the existence of extra conserved quantities is
provided by Kerr-Newmann and Taub-NUT geometry. For the geodesic
motion in the Taub-NUT space, the conserved vector analogous to
the Runge-Lenz vector of the Kepler type problem is quadratic in
$4$-velocities, its components are Killing tensors and they can be
expressed as symmetrized products of Killing-Yano tensors
\cite{Ya, Gibbons, GiRu1, GiRu2, VaVi1, vH}.
It has been shown that in transverse asymptotically flat
space-times, a new set of gravitational charges Y-ADM based on
Killing-Yano tensors are found \cite{KT}. Also, under some
restrictions, pp-wave metrics and Siklos space-times admit
non-generic supercharges \cite{baleanu-baskal}. On the other hand,
there is a relation \cite{Ri1,D1,D2}, called geometric duality,
between spaces admitting irreducible Killing tensors of rank
two and the spaces whose metrics are specified through those
Killing tensors. Further generalizations of Killing tensors and
their existence criteria were discussed \cite{hall, how, HC1,
HC2}. Killing-Yano tensors and their corresponding Killing tensors
have been studied extensively \cite{Carter,CMcL,Visi1,Visi2,Visi3,
Klishevich} in the related context of finding solutions of the
Dirac-equation in non-trivial curved space-time. Moreover, in the
context of generalized Dirac-type operators \cite{tanimoto,
cariglia} the Killing-Yano tensors are indispensable tools.

The aim of this paper is to show that irreducible Killing tensors
can be constructed from third rank Killing-Yano tensors. On the
other hand, we show that the methods, considered until now as
independent, used in the construction of Killing tensors, from
conformal Killing vectors, (known in literature as {\it Koutras
algorithm} \cite{Koutras}, and generalizations \cite{Am-Mah,
ERB}), and the method consisting in the construction of Killing
tensors from Killing-Yano tensors, are interrelated by Hodge
duality. This is true at least in the case when the Hodge dual of
a third rang Killing-Yano tensor is a conformal Killing vector of
gradient type.
In Section 2 some known basic definitions concerning Killing and
Killing-Yano tensors as well as their interrelation are presented.
In Section 3, we give an alternative form for Killing tensors and
Killing tensors equation in terms of Hodge dual of third rank
Killing-Yano tensors making the connection with Koutras algorithm.
In Section 4 we exemplify our results obtained previously on the
Kimura IIC metric. In the first part Killing tensors are to be
expressed in terms of Killing-Yano tensors of various rank and in
the last part Killing tensors are to be constructed from conformal
Killing vectors obtained by Hodge duality from third rank
Killing-Yano tensors.

\section{Killing and Killing-Yano tensors}
In general relativity two generalizations of the Killing vector
equation have been discussed extensively since the discovery of
the Runge-Lenz type vector in the Kerr space-time:

({\it a}) An antisymmetric tensor $f_{\mu_1 \mu_2... \mu_n}$ is
called a Killing-Yano tensor of rank $n$ if the following equation
is satisfied
\begin{equation}\label{ky}
f_{\mu_1 \mu_2... (\mu_n; \lambda)}=0.
\end{equation}

({\it b}) A symmetric tensor field $K_{\mu_1 \mu_2... \mu_n}$ is
said to be a Killing tensor of rank $n$ if:
\begin{equation}\label{k}
K_{(\mu_1 \mu_2...\mu_n; \lambda)}=0.
\end{equation}
From (\ref{ky}) the $n-1$ form field $f_{\mu_1 \mu_2...\mu_{n-1}
\nu}p^\nu$ is parallel transported along affine geodesics with
tangent field $p^\nu$ and from (\ref{k}) we observe that $K_{\mu_1
\mu_2... \mu_n}p^{\mu_1 \mu_2... \mu_n}$ is a first integral of
the geodesic equation.

These two generalizations can be interconnected, constructing a
tensor of rank two as a kind of "square" of Killing-Yano tensor
$f_{\mu_1 \mu_2... \mu_n}$
$$
K_{\mu \nu}=f_{\mu \mu_1 \mu_2... \mu_n}f^{\mu_1 \mu_2...
\mu_n}_{\ \ \ \ \ \ \ \ \ \nu}.
$$
which is a Killing tensor because is symmetric and satisfies the
Killing tensor equation (\ref{k}) with quantity $K_{\mu
\nu}p^{\mu}p^{\nu}=f_{\mu \mu_1 \mu_2... \mu_n}f^{\mu_1 \mu_2...
\mu_n}_{\ \ \ \ \ \ \ \ \ \nu}p^{\mu}p^{\nu}$ the quadratic first
integral generated by Killing tensor. In general, the Killing-Yano
equation (\ref{ky}) has many linear independent solutions
$f^{(i)}_{\mu \mu_1 \mu_2... \mu_n}$, indexed in what follows by
upper latin indices, and a suitable construction  of Killing
tensor $K^{(ij)}_{\mu \nu}$ in this case, is given by
$$
K^{(ij)}_{\mu \nu}=f^{(i)}_{ \mu_1 \mu_2... \mu_{n-1} (\mu}f^{(j)
\mu_1 \mu_2... \mu_{n-1}}_{\nu)}.
$$
Observing that the metric tensor $g_{\mu \nu}$, as well as all
symmetrized products of any Killing vectors $\xi_I$, and in
general, a linear combination of them with constant coefficients,
are Killing tensors, i.e.
\begin{equation}\label{f0}
K_{\mu \nu}=a_0g_{\mu
\nu}+\sum\limits_{I=1}^N\sum\limits_{J=I}^Na_{IJ}\xi_{I(\mu}\xi_{|J|\nu)}
\end{equation}
we can say that all such Killing tensors are called $reducible$
and all other are called $ireducible$.

Briefly speaking, the rank one and four Killing-Yano tensors can
be considered as two trivial cases. In the first case, the
Killing-Yano tensor $f^{(i)}_{\mu}$ is a Killing vector and the
associated Killing tensor
$$
K^{(ij)}_{\mu \nu}=f^{(i)}_{(\mu}f^{(j)}_{\nu)}
$$
is obviously reducible. In the second case, the Killing-Yano
tensor $f^{(i)}_{\mu \nu \lambda \rho}$, solutions of the
equations $f^{(i)}_{\mu \nu \lambda (\rho; \theta)}=0$, is
proportional to the alternating pseudotensor $\eta_{\mu \nu
\lambda \rho}$. In this case the associated  Killing tensor
$$
K^{(ij)}_{\mu\nu}=f^{(i)}_{\alpha\lambda\rho(\mu}f_{\nu)}^{(j)
\alpha\lambda\rho}
$$
is equal with the metric up to a constant.

The Killing-Yano tensors $f^{(i)}_{\mu \nu}$ of rank two are
objects widely discussed in the literature, being indispensable
tools used in the construction of new first integrals of the
geodesic motion, Dirac type operators and gravitational anomalies
investigations. In this case we have twenty independent equations
which have to be solved for six independent components. The
associated Killing tensors are given by
\begin{equation}\label{kt2}
K^{(ij)}_{\mu \nu}=f^{(i)}_{\alpha (\mu}f_{\nu)}^{(j)\alpha}
\end{equation}
some of which are expected to be irreducible.

In order to determine the Killing-Yano tensors
$f^{(i)}_{\mu\nu\lambda}$ of rank three a system with fifteen
independent equations for four independent components
$f^{(i)}_{\mu \nu (\lambda; \rho)}=0$ must to be solved. The
associated Killing tensors are given by
\begin{equation}\label{kt3}
K^{(ij)}_{\mu \nu}=f^{(i)}_{\alpha\beta(\mu
}f_{\nu)}^{(j)\alpha\beta}
\end{equation}
checking which of these are irreducible by comparison with the
definition of reducibility.

\section{Conformal Killing vectors from third rank Killing-Yano tensors}

In this section we extend some of the results obtained in
\cite{Di-Ru} for third rank Killing-Yano tensors and conformal
Killing vectors.  It is well known that if
$f^{(i)}_{\mu\nu\lambda}$ is a third rank Killing-Yano tensor then
the dual reads
\begin{equation}\label{dual}
f^{\ast (i) \mu}=\frac{1}{6}\eta ^{\mu \alpha \beta
\rho}f^{(i)}_{\alpha \beta \rho}.
\end{equation}
Contracting this relation with $\eta_{\mu \alpha \beta \rho}$ we
obtain
$$
f^{(i)}_{\alpha \beta \rho}=-\frac{1}{4}f^{\ast (i) \mu} \eta_{\mu
\alpha \beta \rho},
$$
which enables us to express the Killing tensor (\ref{kt3}) in
terms of $f^{\ast (i) \mu}$ as follows
\begin{equation}\label{Kill-2}
K^{(ij)}_{\mu \nu}= f^{\ast (i)}_{(\mu} f^{\ast (j)}_{\nu)} -
f^{\ast (i)}_{\alpha} f^{\ast (j)\alpha} g_{\mu\nu}.
\end{equation}
with
$$
K^{(ij)}_{(\mu \nu ; \lambda)}= \vartheta^{(j)} f^{\ast
(i)}_{(\mu}g^{}_{\nu \lambda)}+ \vartheta ^{(i)} f^{\ast
(j)}_{(\mu} g_{\nu \lambda)} - (f^{\ast (i)}_{\alpha} f^{\ast
(j)\alpha}_{;(\mu} g_{\nu \lambda)} + f^{\ast (j) \alpha} f^{\ast
(i)}_{\alpha ;(\mu} g_{\nu \lambda)}).
$$
where $\vartheta ^{(i)}$is the conformal factor. The Killing
equation $K^{(ij)}_{(\mu \nu ; \lambda)}=0$ will be satisfied if
$$
 \vartheta^{(j)} f^{\ast
(i)}_{\mu}+ \vartheta ^{(i)} f^{\ast (j)}_{\mu}= f^{\ast
(i)}_{\alpha} f^{\ast (j)\alpha}_{;\mu} + f^{\ast (j) \alpha}
f^{\ast (i)}_{\alpha ;\mu}
$$
or, in another form
$$
f^{\ast (i) \alpha} f^{\ast (j)}_{[\alpha ;\mu ]} + f^{\ast (j)
\alpha} f^{\ast (i)}_{[\alpha ;\mu]}=0.
$$

$\bullet$ {\it Case $(i=j):$ }

This case was partially investigated in \cite{Di-Ru} where was
showed that if $f^{\ast  \mu}$ is null, then the associated
Killing tensor reduces to $K_{\mu \nu}= f^{\ast }_{\mu} f^{\ast
}_{\nu}$ with $f^{\ast  \mu}$ a covariantly constant null Killing
vector , which implies that Killing-Yano tensor
$f_{\mu\nu\lambda}$ is covariant constant. On the other hand, if
$f^{\ast  \mu}$ is non-null, after some algebra, was showed that
$f^{\ast  \mu}$ is a conformal Killing vector $f^{\ast }_{( \mu
;\nu)}=2 \vartheta
 g_{\mu \nu}$ with conformal factor $\vartheta $ and a
gradient field, such that $f^{\ast }_{ \mu ;\nu}=\vartheta g_{\mu
\nu}$. The associated Killing tensor (\ref{Kill-2}) is in this
case given by
\begin{equation}\label{cvcv}
K_{\mu \nu}= f^{\ast}_{ \mu} f^{\ast }_{\nu} - f^{\ast 2}
g_{\mu\nu},
\end{equation}
where $f^{\ast 2}=f^{\ast  }_{\alpha} f^{\ast \alpha}$. If we
define $Q_{\mu \nu} \equiv f^{\ast}_{ \mu} f^{\ast }_{\nu}$, it is
easy to see that $ Q_{\mu \nu} $ is a conformal Killing tensor
$Q_{(\mu \nu ; \lambda)}=g_{(\mu \nu}p_{\lambda)}$ of gradient
type having the associated vector $p_{\lambda}=2 \vartheta
f^{\ast}_{\lambda}$, with $p_{,\lambda}=p_{\lambda}$, allowing to
write for (\ref{cvcv}) the following form
\begin{equation}\label{bar}
K_{\mu \nu}= Q_{\mu \nu}-p g_{\mu\nu}.
\end{equation}
We note that relation (\ref{cvcv}) is also valid for homothetic
($\vartheta=$ const.) Killing vectors of gradient type.

$\bullet$ {\it Case $(i \neq j):$}

In this case, two situations are to be considered:

{\it (a)} First,
$f^{\ast(i) \mu} \equiv  \xi^{ \mu}$ is a null Killing vector, and
$f^{\ast(j) \mu} \equiv f^{\ast \mu}$ a conformal Killing vector,
both of gradient type. The associated Killing tensor
(\ref{Kill-2}) take the form
$$
K_{\mu \nu}= \xi_{ (\mu} f^{\ast}_{\nu)} - \xi_{\alpha} f^{\ast
\alpha} g_{\mu\nu}.
$$
In this case, $ Q_{\mu \nu} \equiv \xi_{ (\mu} f^{\ast}_{\nu)} $
is a conformal Killing tensor $Q_{(\mu \nu ; \lambda)}=g_{(\mu
\nu}q_{\lambda)}$ of gradient type having the associated vector
$q_{\lambda}= \vartheta \xi_{\lambda}$ with $q_{,\lambda}=
q_{\lambda}$. The associated Killing tensor will take the form
$$
K_{\mu \nu}=Q_{\mu \nu}-qg_{\mu\nu}.
$$

{\it (b)} Secondly, in (\ref{Kill-2}) we consider $f^{\ast(i)
\mu}$ and $f^{\ast(j) \mu}$ non-null conformal Killing vectors of
gradient type.  The tensor $Q^{(ij)}_{\mu \nu} $ defined by the
symmetrised product $ Q^{(ij)}_{\mu \nu} \equiv f^{\ast
(i)}_{(\mu} f^{\ast (j)}_{\nu)} $ is a conformal Killing tensor
$Q^{(ij)}_{\mu \nu}=g_{(\mu \nu}\epsilon_{\lambda)}$ of gradient
type having the associated vector $\epsilon_{\mu} = \vartheta
^{(j)}f^{\ast (i)}_{\mu}+\vartheta ^{(i)}f^{\ast (j)}_{\mu}$  with
$\epsilon_{,\mu}=\epsilon_{\mu}$ in which case (\ref{Kill-2}) can
be written
\begin{equation}\label{cvcv2}
K^{(ij)}_{\mu \nu}=Q^{(ij)}_{\mu \nu}-\epsilon g_{\mu \nu}.
\end{equation}
We note that the same discussion can be made in all cases if one
or both Killing vectors $f^{\ast(i) \mu}$, $f^{\ast(j) \mu}$ are
homothetic of gradient type, in which case (\ref{Kill-2}) will
take the appropriate form.

\section{Killing Tensors for Kimura IIC Metric }
Hauser and Malhiot \cite{Hauser-Malhiot} investigated static
spherically symmetric space-times and found all metrics for which
a irreducible Killing tensor of rank two exists obtaining
quadratic first integrals under the assumption that these are
independent of time. Without imposing such assumption Kimura
\cite{Kimura1} generalize the results obtained before by Hauser
and Malhiot and found several metrics having quadratic first
integrals. One of them is know in literature as the Kimura IIC
metric. This metric is of Petrov type $D$, with non-zero
energy-momentum tensor, given by
$$
ds^2=\frac{r^2}{b}dt^2-\frac1{r^2b^2}dr^2-r^2(d\theta^2+\sin^2\theta
d\phi^2)
$$
From Killing vector equations
$$
\xi_{(\mu;\nu)}=0,
$$
we find four Killing vectors
${\bf \xi}^{(i)}=\xi_{(i)}^{\mu}\partial_{\mu}$ given by
$$
\begin{array}{llll}
{\bf \xi}^{(1)}=\frac{\partial}{\partial t} & {\bf \xi}^{(2)}=\cos \phi \frac{\partial}{\partial \theta}+\cot \theta
\sin \phi \frac{\partial}{\partial \phi}\\
{\bf \xi}^{(3)}=-\sin \phi \frac{\partial}{\partial \theta}-\cot
\theta \cos \phi \frac{\partial}{\partial \phi} & {\bf \xi}^{(4)}=\frac{\partial}{\partial \phi}
\end{array}
$$
In covariant form the Killing vectors components reads
$$
\begin{array}{llll}
{\bf\xi}_{(1)}=(\frac{r^2}{b}, 0, 0, 0)&{\bf\xi}_{(2)}=(0, 0, r^2\cos \phi, -r^2 \sin \theta \cos
\theta \sin \phi)\\
{\bf\xi}_{(3)}=(0, 0, r^2\sin \phi, r^2 \sin \theta \cos
\theta \cos \phi)&{\bf\xi}_{(4)}=(0, 0, 0, r^2\sin^2 \theta)
\end{array}
$$

\subsection{Kiling tensors from third rank Killing-Yano tensors}

In what follows we construct Killing tensors from Killing-Yano
tensors of various rank and using (\ref{f0}) we check which of
them are irreducible.

For Killing-Yano tensors of rank two the solution of
corresponding equations (\ref{kt2}) is

$$
f_{23}=r^3\sin\theta,
$$
which exist in any spherically symmetric static space-time
\cite{how, HC2}. The corresponding components of the Killing
tensor are
$$
K_{22}=-r^4,\quad K_{33}=-r^4\sin^2\theta.
$$
It is easy to observe that this tensor can be written as a linear
combination of symmetrized products of Killing vectors
$$
K_{\mu\nu}=-\sum_{i=1}^{3}\xi_{i(\mu}\xi_{|i|\nu)},
$$
identified with the square of the angular momentum, which means
that this Killing tensor is reducible.

By straightforward calculation we obtain the following general
solution for Killing-Yano tensors of rank three,

$$
f_{023}=C_2btr^4\sin\theta+C_1r^4\sin\theta,\quad f_{123}=C_2r\sin\theta
$$
all other components being zero. Therefore, we have two ($i=1,2$),
 independent solutions $f^{(i)}_{\mu\nu\lambda}$, given by,
\begin{equation}\label{sol1}
f^{(1)}_{023}=r^4\sin\theta
\end{equation}
and
\begin{equation}\label{sol2}
f^{(2)}_{023}=btr^4\sin\theta,\quad f^{(2)}_{123}=r\sin\theta
\end{equation}
where we chose the integration constants equals with unity.

Taking into account (\ref{kt3}) we can construct the following
Killing tensors from the Killing-Yano tensors (\ref{sol1}) and
(\ref{sol2}).

$\bullet$ {\it Case 1}: If we consider only the solution (\ref{sol1}) we
obtain for Killing tensor $K^{(11)}_{\mu\nu}$ the following
non-null components
\begin{equation}\label{corrKcase1}
K^{(11)}_{00}=2r^4,\quad K^{(11)}_{22}=-2br^4,\quad K^{(11)}_{33}=-2br^4\sin^2\theta
\end{equation}
which is reducible since can be written as a linear combination of
symmetrized products of Killing vectors
$$
K^{(11)}_{\mu\nu}=2b^2\xi_{0(\mu}\xi_{|0|\nu)}+2b\sum_{i=1}^{3}\xi_{i(\mu}\xi_{|i|\nu)}.
$$

$\bullet$ {\it Case 2}: If we consider only the solution (\ref{sol2}) we
obtain for Killing tensor $K^{(22)}_{\mu\nu}$ the following
non-null components
$$
K^{(22)}_{00}=2b^2t^2r^4,\quad K^{(22)}_{01}=2btr,\quad K^{(22)}_{11}=\frac{2}{r^2},
$$
\begin{equation}\label{corrKcase2}
K^{(22)}_{22}=2b^2r^2(1-bt^2r^2),\quad K^{(22)}_{33}=2b^2r^2(1-bt^2r^2)\sin^2\theta.
\end{equation}
If we define a new Killing tensor $K^{\prime(22)}_{\mu\nu}\equiv
K^{(22)}_{\mu\nu}+ag_{\mu\nu}$, for $a=2b^2$, we obtain the
components
$$
K^{\prime(22)}_{00}=2br^2(bt^2r^2+1),\quad K^{\prime(22)}_{01}=2btr
$$
\begin{equation}\label{corrKpr}
K^{\prime(22)}_{22}=-2b^3t^2r^4,\quad K^{\prime(22)}_{33}=-2b^3t^2r^4\sin^2\theta
\end{equation}
which is irreducible since it is clearly impossible to obtain from
metric and Killing vectors terms which are explicit functions of
$t$.

$\bullet$ {\it Case 3}: Considering both solution (\ref{sol1}) and
(\ref{sol2}) of the third rank Killing-Yano equation, we obtain
the following Killing tensor, with components

\begin{equation}\label{mixt}
K^{(12)}_{00}=2btr^4,\quad K^{(12)}_{01}=r,\quad K^{(12)}_{22}=-2b^2tr^4,\quad K^{(12)}_{33}=-2b^2tr^4\sin^2\theta
\end{equation}
We observe that this tensor can not be obtained as linear
combination of the metric and symmetrized products of Killing
vectors and is also irreducible.

For Killing-Yano tensors of rank four the solution of
corresponding equation (\ref{ky}) yields a system of 4 equations
with one $(f_{0123})$ independent component
$$
f_{0123}=r^2\sin\theta.
$$
We observe that for a four rank Killing-Yano tensor, which is
antisymmetric in all indices, the Killing-Yano equations are
equivalent with $f_{0123;\mu}=0$, and the corresponding Killing
tensor is reducible, proportional with the metric tensor.

\subsection{Killing tensors from conformal Killing vectors}
In this subsection, for Kimura IIC metric, Killing tensors are
constructed from conformal Killing vectors which can be obtained
as dual of third rank Killing-Yano tensors.
\\
From (\ref{sol1}) and (\ref{dual}) we obtain the following
conformal Killing vector
$$
{\bf f}^{\ast (1)}=b\sqrt {b} r^2 \frac{\partial}{\partial r}
$$
with associated conformal factor $\vartheta ^{(1)} = b \sqrt{b}r$.
We observe that ${\bf f}^{\ast}_{ (1)}$ is of gradient type, given
by ${\bf f}^{\ast}_{(1) \lambda}=(-r/{\sqrt {b}})_{,\lambda}$.
From (\ref{cvcv})  with $f^{\ast 2}=-br^2$ we obtain for the
associated Killing tensor components $K_{\mu \nu}$, the same
expressions as in (\ref{corrKcase1}). The same results can be
obtained using (\ref{bar}), with $p_{,\lambda}=\vartheta^{(1)}{\bf
f}^{\ast}_{(1) \lambda }=(-br^2)_{,\lambda}$.
\\
From (\ref{sol2}) and (\ref{dual}) we obtain the following
conformal Killing vector
$$
{\bf f}^{\ast (2)}=-\frac{b \sqrt {b}}{r} \frac{\partial}{\partial
t} + b^2\sqrt {b}tr^2 \frac{\partial}{\partial r}
$$
with associated conformal factor $\vartheta ^{(1)} = b^2
\sqrt{b}tr$. Also, ${\bf f}^{\ast}_{ (2)}$ is of gradient
type, given by ${\bf f}^{\ast}_{(2) \lambda}= (-\sqrt {b}tr)
_{,\lambda}$. In this case, using (\ref{cvcv}) with $f^{\ast
2}=b^2(1-bt^2r^2)$, we obtain for Killing tensor $K_{\mu \nu}$
components the same expressions as in (\ref{corrKcase2}). The
final form (\ref{corrKpr}) is obtained by adding $b^2g_{\mu \nu}$.
On the other hand, (\ref{corrKpr}) can be obtained directly by
using (\ref{bar}), with $p_{,\lambda}=\vartheta^{(2)}{\bf
f}^{\ast}_{(2) \lambda }=(-b^3t^2r^2)_{,\lambda}$.
\\
For the mixt case, from (\ref{Kill-2}) with $f^{\ast (1)}_{\mu }$
given by (\ref{sol1}) and $f^{\ast (2)}_{ \mu }$ given by
(\ref{sol2}) were $f^{\ast (1)}_{\alpha} f^{\ast (2)
\alpha}=-b^2tr^2$, we obtain for Killing tensor components the
same expressions as in (\ref{mixt}). The same results can be
obtained if we use (\ref{cvcv2}) were $\epsilon_{,\lambda} =
\vartheta^{(1)}{\bf f}^{\ast}_{(2) \lambda } +
\vartheta^{(1)}{\bf f}^{\ast}_{(1)
\lambda}=(-b^2tr^2)_{,\lambda}$.

\section{Conclusions}
In this paper, for Kimura IIC metric, higher rank Killing-Yano
tensors were investigated. We have shown that even is impossible
to construct irreducible Killing tensors from two rank
Killing-Yano tensors as in the usual manner, they can be
constructed from third rank Killing-Yano tensors. We obtain, up to
some constants, all Killing tensors known in the literature for
Kimura IIC metric, one reducible and two irreducible. As to our
knowledge the Kimura IIC metric it is the only one metric where
irreducible Killing tensors are constructed from third rank
Killing-Yano tensors.

Therefore, higher rank Killing-Yano tensors prove to be useful
objects for a description of Dirac type operators on a given
background.

\section*{Acknowledgments}
The authors are grateful to Mihai Visinescu for valuable
suggestions and discussions. This work was supported by M.E.C through
AEROSPATIAL-NESSPIAC 105/12.10.2004 and CEEX-CATCG 05-D11-49/07.10.2005 projects.

\end{document}